OpenStudy (anonymous):

find the area of a rectangle tennis court if the perimeter is 480 feet. the length of the court is 6 feet more than twice the width

OpenStudy (anonymous):

the width be 'x' and length be '2x+6' so, \[p = 2(l+b)\] \[2(x+2x+6)=480\]

OpenStudy (anonymous):

\[3x+6=480/2\]

OpenStudy (anonymous):

i dont get this

OpenStudy (anonymous):

the answer is in the thousands

OpenStudy (anonymous):

\[3x=240-6\] \[x=234/3\] \[x=78\] so,width=78 length=2(78)+6=162

OpenStudy (anonymous):

\[area of rectangle = l*w\]

OpenStudy (anonymous):

just plug the values,thats all

OpenStudy (anonymous):

nvm skip this one

OpenStudy (anonymous):

whats your doubt ?

OpenStudy (anonymous):

jus givin me a headache

OpenStudy (anonymous):

do you know perimeter of a rectangle formula ?